### v1

# Correlated interlayer quantum Hall state in alternating twisted trilayer graphene

Abstract
Trilayer graphene offers systematic control of its electronic structure through stacking sequence and twist geometry, providing a versatile platform for correlated states. Here we report magnetotransport in alternating twisted trilayer graphene with a twist angle of about $5^{\circ}$. The data reveal an electron–hole asymmetry that can be captured by introducing layer-dependent potential shifts. At charge neutrality ($\nu_{\rm tot}=0$), three low-resistance states appear, which Hartree–Fock mean-field analysis attributes to emerging spin-resolved helical edge modes resembling those of quantum spin Hall insulators. At $\nu_{\rm tot}=-1$, we also observe suppressed resistance when the middle and bottom layers are each half filled while the top layer remains inert at $\nu=-2$, consistent with an interlayer excitonic quantum Hall state. These results demonstrate correlated interlayer quantum Hall phases in alternating twisted trilayer graphene, encompassing spin-resolved edge transport and excitonic order.
I Introduction
[…]
Multilayer graphene exhibits stacking-dependent electronic structures and transport. In the trilayer case, Bernal (ABA) stacking preserves mirror symmetry about the middle layer, yielding coexisting monolayer- and bilayer-like bands with semimetallic overlap [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Rhombohedral (ABC) stacking instead produces cubic dispersion with an enhanced low-energy density of states, where a perpendicular displacement field polarizes the outer layers and opens a sizable gap [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Beyond these natural stackings, introducing a relative twist between layers reconstructs the electronic spectrum and enables a new class of correlated and topological phases. In the low-angle regime, including magic-angle bilayer, trilayer, and quadrilayer graphene, flat bands host correlated insulating phases, superconductivity, anomalous Hall states, and other symmetry-broken orders [22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Such sensitivity to stacking and twist provides the foundation for exploring correlated phases.
Once the twist angle exceeds the magic angle, the increasing interlayer misalignment produces a substantial momentum mismatch that effectively decouples the electron systems, while interlayer interactions remain significant owing to the atomic interlayer separation. This makes it an ideal ground to realize quantum Hall systems with interlayer excitonic condensation, yet the layers themselves remain independently gate-tunable. For instance, twisted bilayer and twisted double bilayer graphene have provided model systems where excitonic quantum Hall states, possessing gaps significantly larger than those in semiconductor analogues, can be stabilized without the need for insulating barriers [32, 33, 34, 35, 36, 37, 38].
Motivated by the progress described above, here we investigate quantum Hall states in alternating twisted trilayer graphene (ATTG) with a large twist angle close to $5^{\circ}$. Although trilayer graphene has been extensively studied [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 27, 28], the quantizing regime at large twist angles remains unexplored, even though the additional layer degree of freedom is expected to greatly enrich the landscape of emerging correlated phases. In the alternating geometry, where the central layer is rotated relative to the two outer sheets, the system combines features of ABA trilayers with the electronic decoupling characteristic of large twist angles. This interplay generates layer-resolved filling sequences that cannot be realized in bilayers. Indeed, in this regime we uncover an asymmetry between electron and hole doping, along with a series of unconventional low-resistance states at charge neutrality. Mean-field Hartree–Fock analysis indicates that these states are insulating phases with helical edge channels and $R_{xy}=0$, bearing strong resemblance to quantum spin Hall states and pointing to a qualitatively new class of correlated transport in quantum Hall regime. Moreover, at $\nu_{\rm tot}=-1$ we observe a distinct resistance minimum when the middle and bottom layers are simultaneously half filled, a configuration naturally interpreted as an excitonic quantum Hall state stabilized by interlayer coherence. Taken together, our results establish ATTG as a new platform where conventional quantum Hall physics intertwines with correlated phenomena that are unique to the trilayer geometry.
II Landau levels and electron–hole asymmetry
Figure 2: Evolution of interlayer helical-like edge states with displacement field. (a) Longitudinal resistance $R_{xx}$ as a function of $\nu_{\rm tot}$ and $D/\varepsilon_{0}$ at $B=9$ T. Three regions of reduced resistance appear along $\nu_{\rm tot}=0$ (arrows). (b) Schematic representation of (a), showing the layer-resolved lowest Landau levels with the same color code in Fig. 1d, and the vertical splittings reflect broken spin and valley degeneracies. Numbers in the colored boxes mark the layer filling factors. The gray-shaded areas are the low-resistance regions at $\nu_{\rm tot}=0$. (c) Hartree–Fock calculations of layer fillings at $\nu_{\rm tot}=0$ as a function of displacement field. Four stable configurations, labeled i–iv, appear as the displacement field is varied. (d) Edge channel configurations corresponding to the four regions in (c). Black vertical arrows denote spin polarization and horizontal arrows indicate chirality.
The schematic of dual-gated ATTG Device D1, shown in Fig. 1a, has a twist angle of about $5^{\circ}$ (results from Device D2 are provided in Supplementary Note 1 and Supplementary Fig. 1 and 2). Figure 1b presents a color map of the longitudinal resistance $R_{xx}$ as a function of the displacement field ($D/\epsilon_{0}$) and the total filling factor ($\nu_{\text{tot}}$) at $B=4$ T. Several phase transitions are visible, and within the range $60<\nu_{\text{tot}}<60$ their positions are not symmetric about $\nu_{\text{tot}}=0$. This electron–hole asymmetry resembles that reported in dual-gated ABA trilayer graphene [2, 3, 5, 7, 6, 8, 10, 11]. In addition, diamond-shaped boundaries extend along the black dashed line in Fig. 1b, with further diamond features appearing symmetrically with respect to this line. To highlight these features, Fig. 1d provides a summary of the observed transitions. In this plot, the thick red, blue, and green curves denote the lowest Landau levels of the top, middle, and bottom layers, respectively, while the thinner traces correspond to higher-index Landau levels. Numbers in black boxes indicate the Landau level index $N$, and the dotted lines in Fig. 1d can be assigned to crossings between the top and bottom layer Landau levels with the same index.
In Fig. 1d, the vertical feature is attributed to the middle graphene layer, similar to earlier observations in chiral twisted trilayer graphene [39]. The screening effect of the outer layers suppresses the gate response of the middle layer, so its filling remains essentially fixed even as the outer layers change. This distinction leads to the vertical features in Fig. 1b being dominated by the middle layer, while the diagonal traces originate from the outer layers. Unlike the ideal ABA trilayer, where monolayer-like and bilayer-like subbands coexist, the alternating-twist device does not show a bilayer-derived parabolic band. In its place, two Dirac-like dispersions originating from the top and bottom layers coexist with the Dirac cone of the middle sheet, so that the spectrum is effectively composed of three Dirac-like bands, as schematically shown in Fig. 1a. (See Supplementary Fig.3(a)) This difference is most likely caused by residual misalignment between the outer graphene layers, which may result from slight twist-angle deviations, interlayer sliding during assembly, or both.
To interpret the experimental results, we performed tight-binding calculations using the potential profile shown in the inset of Fig. 1a, based on the Slonczewski–Weiss–McClure parameters for ABA trilayer graphene [4, 6, 9, 10, 11]. In this model the displacement field $D$ induces a potential difference $\Delta$ between the outer layers, while an additional parameter $\delta$ accounts for the residual charge imbalance between the outer and middle sheets. For $D=0$, only $\delta$ is present, giving $V_{\rm top}=V_{\rm bot}=+\delta$ and $V_{\rm mid}=-2\delta$. With finite $D$, the layer potentials become
$V_{\rm top}=+\tfrac{\Delta}{2}+\delta,\quad V_{\rm mid}=-2\delta,\quad V_{\rm bot}=-\tfrac{\Delta}{2}+\delta.$
A small imbalance of order a few meV is typically required even in ABA trilayers, so we set $\delta=8$ meV [4, 6, 9, 10]. With this choice the calculated density of states map in Fig. 1c reproduces the main features of the experiment. Circles highlight crossings of the lowest Landau levels, whose positions coincide with the transition points observed in the resistance map of Fig. 1b and summarized in Fig. 1d. The crossing of the first excited Landau levels predicted by the calculation also matches the corresponding transition in the data. This agreement shows that the potential model captures the essential electronic structure of the alternating trilayer. Further results at higher magnetic fields are presented in Supplementary Note 2 and Supplementary Fig. 3.
III Insulators with Helical Edge Transport at $\nu_{\rm tot}=0$
At higher magnetic fields the effects of particle–hole asymmetry and charge imbalance become more pronounced. Figure 2a shows the longitudinal resistance in the lowest Landau level at $B=9$ T and $T=1.5$ K. Some fractional quantum Hall states are visible in the lowest Landau levels of the individual layers, and crossings between symmetry-broken states of different layers appear more clearly than at lower fields. The most striking feature occurs at $\nu_{\rm tot}=0$, where the resistance remains largely high across most displacement field, but three distinct regions of reduced resistance emerge, highlighted by arrows in Fig. 2a.
The schematic in Fig. 2b extends the picture of Fig. 1d, making explicit the symmetry-broken states that emerge under displacement-field tuning. Green, blue, and red lines denote the Landau levels of the top, middle, and bottom layers, respectively, and their vertical splittings reflect the lifting of spin and valley degeneracies. Numbers in the colored boxes indicate the layer filling factors, and the three shaded regions correspond to $(\nu_{\rm top},\nu_{\rm mid},\nu_{\rm bot})=(-2,1,1)$, (1,2,1)(-1,2,-1), and (1,1,2)(1,1,-2) for positive, near-zero, and negative displacement fields. These configurations match the positions of the resistance minima at $\nu_{\rm tot}=0$ in Fig. 2a, showing that the low-resistance states originate from specific combinations of symmetry-broken Landau levels across the three layers.
These symmetry-breaking patterns are further supported by our mean-field Hartree–Fock calculations at $\nu_{\rm tot}=0$, which track the layer-resolved filling factors as a function of displacement field, as shown in Fig. 2c. In this calculation, we model each layer as continuum Dirac fermions subject to a uniform magnetic field, with the asymmetry $\delta$ determined from our tight-binding analysis (Supplementary Note 3). The three colored traces give $\nu_{\rm top}$, $\nu_{\rm mid}$, and $\nu_{\rm bot}$, whose sum is constrained to zero at all fields. Four different configurations, labeled i–iv, appear as the displacement field is swept. In direct comparison with the data in Fig. 2a, states ii and iv coincide with high-resistance sectors, while states i and iii align with the low-resistance regions at $\nu_{\rm tot}=0$.
The underlying edge-channel configurations of these correlated insulators are illustrated in Fig. 2d. Left- (right-) pointing arrows indicate electron-like (hole-like) chirality for each layer-resolved edge mode. In the high-resistance cases (ii and iv), the relevant fillings include $\pm 2$ on adjacent layers, which generate counter-propagating channels of the same spin; interlayer scattering is then efficient and $R_{xx}$ is enhanced [40, 41]. In contrast, in configurations i and iii the numbers of right- and left-moving modes for a fixed spin do not match. For instance, in case i the left-moving spin-$\uparrow$ mode has no same-spin partner for scattering. By contrast, the two right-moving and one left-moving spin-$\downarrow$ modes are expected to recombine into an effective single right-moving spin-$\downarrow$ mode. Because spin-$z$ is approximately conserved in graphene, inter-spin scattering is forbidden, leaving the two modes—one left-moving spin-$\uparrow$ and one effective right-moving spin-$\downarrow$—largely intact. This suppression of edge backscattering leads to a reduced $R_{xx}$. This state is marked by $R_{xy}=0$ and spin-resolved helical edge modes, which are the defining transport signatures of quantum spin Hall insulators [40, 41], though realized here under broken time-reversal symmetry induced by the magnetic field.
Such spin-resolved interlayer helical edge configurations are a unique feature of ATTG, realized for example by the filling sequence $(\nu_{\rm top},\nu_{\rm mid},\nu_{\rm bot})=(\pm 1,\pm 1,\mp 2)$. When $\nu_{\rm tot}=0$ the numbers of electron- and hole-like channels compensate, so counter-propagating modes with opposite spin can form a robust helical state, owing to the conservation of spin-$z$ in graphene under magnetic fields. For $\nu_{\rm tot}\neq 0$ this compensation is lost, and an extra co-propagating channel provides an equilibration pathway that makes such pairs fragile. This explains why analogous states have not been observed in bilayers: in twisted bilayer graphene the (1,+1)(-1,+1) configuration at $\nu_{\rm tot}=0$ can host a state [40], but the (2,+1)(-2,+1) configuration lacks a partner channel and reduces to a conventional quantum Hall state. In ATTG the additional $\nu=+1$ layer supplies this partner, yielding mixed helical and non-helical edge modes that remain visible.
To further confirm this interpretation, we carried out non-local transport measurements, which directly test the helical character of the edge channels. The details of these measurements, including probe configurations and field dependence, are provided in Supplementary Note 4, Supplementary Fig. 4, and Supplementary Table 2.
IV Field dependence of helical edge transport
Figure 3: Magnetic field dependence of helical-like edge states. (a) Contour map of longitudinal resistance $R_{xx}$ as a function of carrier density and displacement field at $B=11$ T. (b) Line traces of $R_{xx}$ as a function of displacement field at $n_{\rm tot}=5\times 10^{10}\,\mathrm{cm}^{-2}$, taken along the green dotted line in (a), for magnetic fields between 12 T and 6 T. (c) Contour map of Hall resistance $R_{xy}$ measured under the same conditions as in (a). (d) Line traces of $R_{xy}$ corresponding to the green dotted line in (c), at $n_{\rm tot}=5\times 10^{10}\,\mathrm{cm}^{-2}$ for fields between 12 T and 6 T.
The magnetic field dependence of the helical edge states is summarized in Fig. 3. At $B=11$ T, the longitudinal resistance map in Fig. 3a shows clear reductions in $R_{xx}$ along $\nu_{\rm tot}=0$, consistent with the features identified in Fig. 2. These reductions appear most strongly in regions composed of two $\pm 1$ and one $\mp 2$ fillings: (-2,1,1), (-1,2,-1), (1,1,-2), indicated by the white arrows. Additional suppression is observed for $(\pm 2,0,\mp 2)$ compared with $(0,\pm 2,\mp 2)$, which can be attributed to the larger spacing between edge channels that reduces interlayer backscattering. Line cuts taken at $n_{\rm tot}=5\times 10^{10}\,\mathrm{cm}^{-2}$, shown in Fig. 3b, demonstrate that the dips in $R_{xx}$ persist from $B=12$ T down to 6 T, but their depth decreases steadily as the field is lowered. This trend indicates that the stability of the low-resistance states relies on the field-induced spin and valley splitting that enters the Hartree–Fock analysis of Fig. 2c.
The Hall resistance exhibits the same evolution. In Fig. 3c, the contour map at 11 T shows reduced $R_{xy}$ in the same displacement field windows where $R_{xx}$ is suppressed. The corresponding line cuts in Fig. 3d reveal that this suppression becomes weaker at lower fields, showing the same trend as the longitudinal response. As the magnetic field decreases, the degeneracy lifting of the $n=0$ Landau level is reduced, interlayer scattering is restored, and the low-resistance state gradually disappears.
V Interlayer excitonic quantum Hall state at $\nu_{\rm tot}=-1$
Figure 4: Interlayer coherence quantum Hall states at $\nu_{\rm tot}=-1$. (a) Longitudinal resistance $R_{xx}$ plotted as a function of $\nu_{\rm tot}$ and $D/\varepsilon_{0}$. (b) Schematic map of (a). Blue and red shading denotes filling of the middle and bottom layers, respectively, while the top layer remains fixed at $\nu_{\rm top}=-2$. The black solid line at $\nu_{\rm tot}=-1$ corresponds to integer quantum Hall states with $(\nu_{\rm mid},\nu_{\rm bot})=(0,1)$ and $(1,0)$. Dark red (blue) lines in the interval $2<\nu_{\rm tot}<-1$ indicate fractional quantum Hall states of the bottom (middle) layer. The black dot marks a low-resistance region. (c) Evolution of $R_{xx}$ with displacement field at $\nu_{\rm tot}=-1$ for several magnetic fields. The arrow highlights resistivity minima. Curves are offset vertically and horizontally for clarity.
Another intriguing state of ATTG emerges at $\nu_{\rm tot}=-1$, where an interlayer excitonic state develops. Figure 4a presents a magnified window of the $(\nu_{\rm tot},D/\varepsilon_{0})$ plane over the range $2<\nu_{\rm tot}<0$. In this regime the top layer is locked at $\nu_{\rm top}=-2$, leaving the middle and bottom sheets to host the relevant lowest Landau level physics. Fractional quantum Hall features at $\nu=-5/3$ and 4/3-4/3 arise from one of these two active layers, as summarized in Fig. 4b.
At $\nu_{\rm tot}=-1$ and $D/\varepsilon_{0}\approx 165$ mV/nm, a narrow dip develops in $R_{xx}$. The schematic in Fig. 4b identifies this point as the balanced configuration $\nu_{\rm mid}=\nu_{\rm bot}=1/2$, corresponding to half filling of the lowest Landau level in both adjacent layers. This is the standard condition for interlayer coherent exciton formation in graphene bilayers, realized here without an insulating barrier between the two active sheets[33, 35, 37, 36, 38]. These observations highlight how a hallmark of bilayer quantum Hall physics, which is the interlayer excitonic state at half filling, acquires a new realization in twisted trilayers, in which one layer remains inert and the remaining two layers condense into an excitonic phase. The observed minima of $R_{xx}$ can therefore be understood as the combined response of a conventional edge channel in the top layer and a coherent electron–hole condensate in the middle–bottom subsystem, showing how conventional quantum Hall order and interlayer correlated phases can intertwine in a trilayer platform enabled by large-angle decoupling.
The field dependence of this feature is shown in Fig. 4c. Line traces of $R_{xx}$ at $\nu_{\rm tot}=-1$ reveal that the dip (black arrows) weakens steadily as the magnetic field is lowered from 9 T to 5 T. This trend is consistent with an interaction-driven coherent state whose stability depends on the exchange-enhanced gap at half filling: as the field decreases, the gap diminishes and the excitonic signature fades [35, 37, 36, 38].
VI Discussion
In conclusion, we have investigated the quantum Hall states of ATTG and identified correlated interlayer quantum Hall phases that are not accessible in conventional bilayer and trilayer systems. The middle layer responds primarily to carrier density while the outer layers are strongly influenced by both density and displacement field, producing a landscape of filling factor combinations unique to this geometry. At $\nu_{\rm tot}=0$, a reduction in resistance appears. Comparison with Hartree–Fock calculations establishes that these low-resistance states arise from spin-resolved helical edge modes, closely analogous to those in a quantum spin Hall insulator. Unlike the bilayer case, here all three graphene layers participate, enabling configurations such as $(\pm 2,0,\mp 2)$ and $(0,\pm 2,\mp 2)$, which reveal how interlayer separation governs the strength of backscattering. This demonstrates that alternating trilayers provide a broader and more tunable platform for exploring spin-polarized edge conduction than bilayer systems.
A further characteristic of this system is the emergence of an interlayer excitonic quantum Hall state at $\nu_{\rm tot}=-1$. In this regime the top sheet forms a conventional integer quantum Hall state at $\nu_{\rm top}=-2$, while the middle and bottom sheets are simultaneously half filled, creating conditions for exciton condensation across adjacent layers. The resulting suppression of $R_{xx}$ mirrors excitonic states reported previously in twisted bilayer and double bilayer graphene, but here it coexists with a fully filled outer layer, providing a new trilayer realization of a bilayer hallmark. The field dependence of this feature confirms its interaction-driven origin. In summary, ATTG combines conventional quantum Hall features with emergent interlayer correlated phases, highlighting correlated interlayer quantum Hall phases unique to trilayer geometries with tunable layer asymmetry.
VII Methods
Device fabrication
We fabricated a graphite/h-BN/twisted trilayer graphene/h-BN/graphite/h-BN (from bottom to top) using dry pick up method with Elvacite stamp. Both top and bottom graphite were used as gates. The assembled heterostructure was patterned into a Hall bar geometry using electron-beam lithography followed by reactive ion etching with mixture of $\text{CF}_{4}$ and $\text{O}_{2}$ at power of 40 W. Finally, a metal electrode was deposited with Cr/Au (5/100nm) by e-beam evaporator.
Transport measurements
Transport measurements were performed with the help of low frequency lock-in techniques using an ac current of 100 nA and a frequency of 13.333 Hz in a He4 refrigerator with a base temperature of 1.5 K. In order to reduce contact resistance in areas not covered by the graphite gates, the silicon substrate was biased at $\pm$70 V during the measurements.
Capacitance model
The total filling factor $\nu_{\text{tot}}$ and the displacement field $D/\epsilon_{0}$ were determined within a simple capacitance model as
$\nu_{\text{tot}}=\frac{n_{\text{tot}}h}{eB}=\frac{(C_{\text{t}}V_{\text{t}}+C_{\text{b}}V_{\text{b}})h}{e^{2}B},\qquad\frac{D}{\epsilon_{0}}=\frac{C_{\text{t}}V_{\text{t}}-C_{\text{b}}V_{\text{b}}}{2\epsilon_{0}}.$
Here $n_{\text{tot}}$ is the carrier density, $C_{\text{t(b)}}$ the capacitance per unit area of the h-BN dielectric between the top (bottom) gate and the trilayer graphene, and $V_{\text{t(b)}}$ the applied top (bottom) gate voltage [33, 35, 38, 39, 41]. $e$ denotes the elementary charge, $h$ Planck’s constant, and $B$ the magnetic field.
Tight-binding calculation
Following our approach for magic-angle alternating twist graphenes [42], we here too model the system within a tight-binding framework in the localized orbital basis $|i\rangle$} of carbon sites [43]:
$\hat{H}=\sum_{i}\epsilon_{i}|i\rangle\langle i|+\sum_{i,j}t_{ij}|i\rangle\langle j|,$ (1)
where $\epsilon_{i}$ are onsite terms from the F2G2 graphene parametrization [44], corrected by the onsite energy shifts $\Delta$ and $\delta$ as discussed in the main text, and $t_{ij}$ are distance-dependent hopping amplitudes. In our SHE model [45], intralayer couplings follow the F2G2 model, while interlayer tunneling is described by the two-center approximation:
$t_{ij}^{\rm inter}=S,\exp\left(\frac{c_{ij}-p}{q}\right)t^{\rm TC}_{ij},$ (2)
with p=3.25p=3.25Å, q=1.34q=1.34Å, and $c_{ij}$ the vertical separation. The two-center term is
$t^{\rm TC}{ij}=V{pp\pi}(r_{ij})\left[1-\left(\frac{c_{ij}}{r_{ij}}\right)^{2}\right]+V_{pp\sigma}(r_{ij})\left(\frac{c_{ij}}{r_{ij}}\right)^{2},$ (3)
where $r_{ij}=|{\bm{r}}_{ij}|$, with
$V_{pp\pi}(r)=V_{pp\pi}^{0}\exp\left(-\frac{r-a_{0}}{r_{0}}\right)$ (4)
and
$V_{pp\sigma}(r)=V_{pp\sigma}^{0}\exp\left(-\frac{r-c_{0}}{r_{0}}\right).$ (5)
We set $a_{0}=1.42$Å, $c_{0}=3.35$Å, $V_{pp\pi}^{0}=-2.7$eV, $V_{pp\sigma}^{0}=0.48$eV, and $r_{0}=0.184a$. Interactions vanish beyond 4.94.9Å in accordance with Ref. [46]. The prefactor $S$ is commonly adjusted to reproduce the first magic-angle near $1.1^{\circ}$[47]. Since our focus is on a large angle of $\sim 5^{\circ}$, where moiré reconstruction is negligible, we adopt the rigid-lattice value $S=0.895$ without further correction. For completeness, we note that energy minimization relaxations, using LAMMPS [48] with REBO2 intralayer[49] and DRIP interlayer [50], fitted to EXX-RPA data [51], produce only minor modifications at such large twist and do not affect the observables considered. For electronic properties, we apply the Lanczos method with random-phase vectors to compute the density of states [52]:
[…]
with broadening parameter $\eta$. Magnetic fields are included via a Peierls substitution of phases in the hopping amplitudes [53].
Mean-field Hartree-Fock calculation
We theoretically model our system of ATTG with the continuum Dirac Hamiltonian consisting of the phenomenological short-range interaction terms [35], potential energy terms induced by the displacement fields and Zeeman effect, and the capacitive energy term modeled by three infinite plate (See Supplementary Note 4 of SI for details). We focus on the 12-dimensional lowest Landau levels, given by layer $\times$ valley $\times$ spin = $3\times 2\times 2$. We determine the ground state using the Hartree–Fock method within a valley-, layer-, and spin-polarized ansatz. Applying Hartree-Fock method and Fierz identity together, we reduce the interactions into a sum of one-body pseudospin operators expressed in the direct product of Pauli-matrix basis, and represent the state by 12-component spinor [54]. We performed a brute-force search restricted to diagonal projectors. Further details are provided in Supplementary Note 4 of the SI.
VIII Data availability
Source data for all main text figures are included in the Supplementary Information. Any other data supporting the findings of this study are available from the corresponding author upon request. Source data are provided with this paper
IX Code availability
Key algorithms of theoretical calculations are fully clarified in the Supplementary Information.